1. A combined gas-steam power cycle uses a single gas turbine cycle for the air cycle and a simple Rankine cycle for the water vapor cycle. Atmospheric air enters the compressor at a rate of 88.2 lbm / s, at 14.7 psia and 59 ° F, and the maximum gas cycle temperature is 1,742 ° F. The pressure ratio in the compressor is 7. The isentropic efficiency of both the compressor and the turbine is 80%. Gas exits the heat exchanger at 392 ° F. The Rankine cycle operates between the pressure limits of 870.4 psia and 1.45 psia. Water enters 752 ° F into the turbine. The isentropic efficiency of both the turbine and the pump is 80%. Use software and obtain the gas turbine outlet temperature, the combined cycle thermal efficiency, the change in entropy in the turbines, and the power generated in the combined cycle.
Combustion chamber Gas turbine Gaske Compressor € Air Heat exchanger Exhaust gases turbine Scam cycle Pump Condenser


cyclepad if you can ;)
Answer #1

1) Gas Turbine outlet temperature:

Given Data:

Mass flow rate m=88.2lb/s=40kg/s

P1=14.7psi=1.013bar

T1=59 F=288K

T3=1742F=1223K

pressure ratio=7

V(gamma)=1.41

n isen=0.8

As per Temperature, Pressure relation

T2/T1=(p2/p1)^(v-1)/v T3/T4=(P4/P3)^(v-1)/v

T2/288=(7)^(0.4)/(1.41) 1223/T4=(7)^(0.4)/(1.41)

T2=500K T4=708K

For turbine,

Isentropic efficiency=Nt=(h3-h4s)/(h3-h4) (h=mcpT)

0.8 =(1223-T4s)/(1223-708)

T4s=811K